Two bumper cars collide head on. Bumper car 1 has a weight of 150 lb and is moving with a speed of 6 ft/s. Bumper car 2 has a weight of 170 lb and is moving with a speed of 4 ft/s. If no energy is lost during the collision, A) what are the velocities of the two cars after the collision? B) What is the magnitude of the average force between the two cars during the collision if it lasts 0.7 seconds?
Answer #1

150 16 110.16 O Before collision 6fts uftis Isolb - Tolb - V2 After collision momentum consevation - 150x6 - 170x4 = 170xV2-Vi= 4.625 LHS velocity after collision V2 = 5.375 HA1S - - iD Impulsive force for cara = may 1 = 150 (6 - (-4.625) 0-7 7 2276